Corn Syrup worth P44.80 a gallon and maple syrup worth P336 a gallon are mixed to form a mixture worth P132.16 a gallon. How many gallons of each kind of syrup are needed to make 50 gallons of the mixture?
Answer by mananth(16946)

---------price ---------------- quantity
corn syrup 44.8 ---------------- x gallons
maple syrup 336 ---------------- 50 - x gallons
Mixture 132.16% ---------------- 50

44.8 x + 336 ( 50 - x ) = 132.16 * 50

44.8x + 16800 - 336 x = 6608
44.8 x - 336 x = 6608 - 16800
-291.2 x = -10192
/ -291.2
x = 35 gallons corn syrup
15 gallons maple syrup